OpenStudy (llivingston2):

What is the remainder when (x3 − 4x2 − 12x + 9) is divided by (x + 2)?

OpenStudy (mhchen):

Do you know long or synthetic division?

OpenStudy (llivingston2):

no

OpenStudy (mhchen):

Okay.......so did the teacher ever show how to do this problem during class? :o I can explain this in 2 ways, and I'm not sure which way your teacher wants it

OpenStudy (mhchen):

I'll use long division:

DivineSolar (divinesolar):

Long division would be correct :P

OpenStudy (mhchen):



DivineSolar (divinesolar):

Use the equation button. :)

OpenStudy (mhchen):

Okay so I did the first step. Do you see how I did it.

OpenStudy (mhchen):

Equation button won't let me do long-division.

OpenStudy (mhchen):

or I'm too noob to know how..

OpenStudy (llivingston2):

I kind of see

OpenStudy (mhchen):

Yeah, it's just doing that process over and over again.

OpenStudy (mhchen):

just like normal division, and whatever is left is the remaindr.

DivineSolar (divinesolar):

Well instead of having you go through a long process lets just have him do this :). Save you some of the work. Subtract -6x^2 - 12x from -6x^2 -12x + 9 to get your remainder.

OpenStudy (llivingston2):

ok......

DivineSolar (divinesolar):

I can give you the entire process in words if need be later on :)

OpenStudy (llivingston2):

yes please @DivineSolar

OpenStudy (llivingston2):

wait so is the remainder 9

DivineSolar (divinesolar):

Divide the leading co-efficients of the numerator \[\frac{ x^3 - 4x^2 - 12x + 9 }{ x + 2 }\] and the divisor aka x + 2 : x^3/x = x^2 aka our quotient. And yes the remainder is 9

OpenStudy (llivingston2):

yay thanks

DivineSolar (divinesolar):

Sure.